OpenStudy (anonymous):

Integrate (2^Ln(1/x)) / (x), with respect to x. @Calculus1

myininaya (myininaya):

\[\int\limits_{}^{}\frac{2^{\ln(\frac{1}{x})}}{x} dx\]?

OpenStudy (anonymous):

yep

myininaya (myininaya):

let u=ln(1/x)=ln(1)-ln(x)=-ln(x) du=-1/x dx

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\int\limits{\frac{2^{\ln(\frac{1}{x})}}{x}}dx\] let u=ln(1/x) du=-1/x dx \[\int\limits{-2^udu} = \frac{-2^u}{\ln(2)}\] \[=\frac{-2^{\ln(\frac{1}{x})}}{\ln2}\]

myininaya (myininaya):

\[\int\limits_{}^{} 2^u (-du)\]